OpenStudy (abmon98):

If we repeated the process of (0!+0!)+0!)!...) will the result reach infinity?

OpenStudy (bubblegum.):

u must add a lotta brackets on left too \(((((...(0!+0!)+0!)...)\) and yeah i think it is infinity

zepdrix (zepdrix):

You're really just doing this, right?\[\large\rm (((...(2!+1)!+1)!+1)!+...\]Since the inner most part is 0!+0! = 2. Ya looks like it grows pretty crazy :) What a mess, heh

OpenStudy (abmon98):

I am getting increasing number of factorials

zepdrix (zepdrix):

Ya, and they increase pretty quickly don't they? :)\[\large\rm (((...(2!+1)!+1)!+1)!+...\]\[\large\rm (((...(3!+1)!+1)!+...\]\[\large\rm (((...(7!+1)!+...\]and the next one is huuuuuge

OpenStudy (abmon98):

then whats the factorial of infinity

OpenStudy (bubblegum.):

infinity

OpenStudy (bubblegum.):

its just-> \(\infty \times (\infty-1) \times (\infty-2)...3 \times 2 \times 1\) what ever u multiply with \(\infty\) you will always get \(\infty\)

OpenStudy (abmon98):

I thought that there is something special about it thanks..

OpenStudy (bubblegum.):

np :) few days ago i was thinking about factorial and thought about \((-1)!\) i entered it on wolfram and i got this- http://www.wolframalpha.com/input/?i=%28-1%29%21

OpenStudy (abmon98):

(-1)!=0!/0=1/0=undefined

OpenStudy (anonymous):

silence